If sample space S={e1,e2,e3,e4}, find P(e1 ) and P( e2) if P(e3)=1/3 and P(e2)=3P(e1)
1
Expert's answer
2021-07-14T16:11:43-0400

"S=\{ e_1,e_2,e_3,e_4\}"


Given,

"P(e_3)=\dfrac{1}{3}\\P(e_2)=3P(e_1)"


Now,

"P(e_1)+P(e_2)+P(e_3)+P(e_4)=1\\\implies P(e_1)+P(e_2)+P(e_4)=1-\dfrac{1}{3}=\dfrac{2}{3}\\\ \\\implies P(e_1)+3P(e_1)+P(e_4)=\dfrac{2}{3}\ [\because P(e_2)=3P(e_1)]\\\ \\\implies 4P(e_1)+P(e_4)=\dfrac{2}{3}"


Since, the value of "P(e_4)" is not given , so we assume the value of "P(e_4)"

Let "P(e_4)=\dfrac{1}{6}"


So,

"\implies 4P(e_1)=\dfrac{2}{3}-\dfrac{1}{6}=\dfrac{3}{6}=\dfrac{1}{2}\\\ \\\implies P(e_1)=\dfrac{1}{8}"


and

"P(e_2)=3P(e_1)\\\implies P(e_2)=\dfrac{3}{8}"